Check List for Particle : Ξ_{b}^{−}

April 19, 2009

Quantity  Us  PDG  EvtGen  Quantity  Us  PDG  EvtGen
PDG  5132  Mass  5792.9

Code

Width 0 Gen Xi_b-

Name

Decay 0.000417 m.

Length

Latex Name Ξ_{b}^{−}  Mass Gen.  Width Gen.  Type baryon
Type  normal  Con. Mass 0  Rating  ∗ ∗ ∗  Variable  fixed

Ratio

Stable  unstable  Spin  1/2  Charge  −1  Colour  0
Lower Cut  0  Upper cut 0  PDG link  s060

• Is the description present and correct?

The mass is taken from [?] . The lifetime is the experimentally measured value from [?].

No branching ratios have been measured. We have therefore chosen to add a number of two body modes with branching ratios calculated from various theoretical papers and model the remaining modes using partonic decays. The ratios of the partonic decays were calculated using the colour factor for the relative rates of the leptonic and direct b → c decays, colour exchanged decays were added with one third of the rate for the direct decays. The rates of the different quarks produced in the weak decays were calculated using the CKM matrix elements.

• References?

Branching Rating  Outgoing  Description  Decayer  EvtGen
Ratio  Particles
0.203900  ∗∗  s, c̄, c, (sd)_{0}  Herwig::WeakPartonic::WeakPartonic100
0.201600  ∗∗  d, ū, c, (sd)_{0}  Herwig::WeakPartonic::WeakPartonic100
0.091300  ∗∗  c, ū, d, (sd)_{0}  Herwig::WeakPartonic::WeakPartonic100
0.089460  ∗∗  c, c̄, s, (sd)_{0}  Herwig::WeakPartonic::WeakPartonic100
0.073200  ∗∗  Ξ_{c}^{0}, μ^{−}, ν̄_{μ}  Herwig::SemiLeptonicBaryon::BaryonHeavy1
0.073200  ∗∗  Ξ_{c}^{0}, e^{−}, ν̄_{e}  Herwig::SemiLeptonicBaryon::BaryonHeavy1
0.043000  ∗∗  Ξ_{c}^{0}, D_{s}^{∗−}  Herwig::BaryonFactorized::BottomBaryonVector

Table 1: The decay modes of the Ξ_{b}^{−} (continues).

Branching Rating  Outgoing  Description  Decayer  EvtGen
Ratio  Particles
0.037000  ∗∗  Ξ_{c}^{0}, a_{1}^{−}(1260)  Herwig::BaryonFactorized::BottomBaryonVector
0.027000  ∗∗  Ξ_{c}^{0}, D_{s}^{−}  Herwig::BaryonFactorized::BottomBaryonScalar
0.026000  ∗∗  Ξ_{c}^{0}, ρ^{−}  Herwig::BaryonFactorized::BottomBaryonVector
0.023000  ∗∗  μ^{−}, ν̄_{μ}, c, (sd)_{0}  Herwig::WeakPartonic::WeakPartonic100
0.023000  ∗∗  e^{−}, ν̄_{e}, c, (sd)_{0}  Herwig::WeakPartonic::WeakPartonic100
0.021100  ∗∗  Ξ_{c}^{0}, τ^{−}, ν̄_{τ}  Herwig::SemiLeptonicBaryon::BaryonHeavy1
0.012300  ∗∗  s, ū, c, (sd)_{0}  Herwig::WeakPartonic::WeakPartonic100
0.012000  ∗∗  d, c̄, c, (sd)_{0}  Herwig::WeakPartonic::WeakPartonic100
0.011000  ∗∗  τ^{−}, ν̄_{τ}, c, (sd)_{0}  Herwig::WeakPartonic::WeakPartonic100
0.009000  ∗∗  Ξ_{c}^{0}, π^{−}  Herwig::BaryonFactorized::BottomBaryonScalar
0.004800  ∗∗  c, c̄, d, (sd)_{0}  Herwig::WeakPartonic::WeakPartonic100
0.004800  ∗∗  c, ū, s, (sd)_{0}  Herwig::WeakPartonic::WeakPartonic100
0.002200  ∗∗  s, c̄, u, (sd)_{0}  Herwig::WeakPartonic::WeakPartonic100
0.002200  ∗∗  d, ū, u, (sd)_{0}  Herwig::WeakPartonic::WeakPartonic100
0.001500  ∗∗  Ξ_{c}^{0}, D^{∗−}(2010)  Herwig::BaryonFactorized::BottomBaryonVector
0.001500  ∗∗  Ξ_{c}^{0}, K^{∗−}  Herwig::BaryonFactorized::BottomBaryonVector
0.001000  ∗∗  Ξ^{−}, J/ψ(1S)  Herwig::BaryonFactorized::BottomBaryonVector
0.001000  ∗∗  Ξ_{c}^{0}, D^{−}  Herwig::BaryonFactorized::BottomBaryonScalar
0.000840  ∗∗  Ξ^{−}, ψ(2S)  Herwig::BaryonFactorized::BottomBaryonVector
0.000800  ∗∗  μ^{−}, ν̄_{μ}, u, (sd)_{0}  Herwig::WeakPartonic::WeakPartonic100
0.000800  ∗∗  e^{−}, ν̄_{e}, u, (sd)_{0}  Herwig::WeakPartonic::WeakPartonic100
0.000700  ∗∗  Ξ_{c}^{0}, K^{−}  Herwig::BaryonFactorized::BottomBaryonScalar
0.000300  ∗∗  τ^{−}, ν̄_{τ}, u, (sd)_{0}  Herwig::WeakPartonic::WeakPartonic100
0.000200  ∗∗  Σ_{c}^{0}, K^{−}  Herwig::Hw64::DecayME0
0.000100  ∗∗  Ξ′_{c}^{0}, π^{−}  Herwig::Hw64::DecayME0
0.000100  ∗∗  s, ū, u, (sd)_{0}  Herwig::WeakPartonic::WeakPartonic100
0.000100  ∗∗  d, c̄, u, (sd)_{0}  Herwig::WeakPartonic::WeakPartonic100

Table 1: The decay modes of the Ξ_{b}^{−}.

Total branching ratio is 1.